# ShortcutsBench: A Large-Scale Real-world Benchmark for API-based Agents

### v2

@@ 5 Conclusion
In this paper, we introduce ShortcutsBench, a benchmark for evaluating API-based agents. To the best of our knowledge, ShortcutsBench is the most realistic, rich, and comprehensive benchmark of its kind. Our findings indicate that for agents built on the most advanced LLMs, the primary bottleneck is API selection. For the most cost-effective LLMs, there is considerable room for improvement in both API selection and parameter filling. Additionally, we identified a significant deficiency in the agents’ awareness of requesting necessary information.
@@ Appendix A Appendix / supplemental material
A.1 Dataset Acquisition Process
In this section, we introduce more details about the dataset acquisition.
Regarding data acquisition, we first use search engines to identify popular public shortcut-sharing sites. We found a total of $14$ sites. These sites include:
@@ A.1 Dataset Acquisition Process
In addition to collecting iCloud Links (URLs) from shortcut-sharing sites, these sites also provide information such as the shortcut’s name (NameInStore), functional description (DescriptionInStore), category (CategoryInStore), number of downloads (Downloads), favorites (Favorites), reads (Reads), and ratings (Rates). Most shortcuts include NameInStore and DescriptionInStore (except for a few obtained from Reddit), while the availability of other fields varies slightly depending on the specific shortcut-sharing site.
After deduplicating based on “iCloud link” (Apple, 2024c), we attempt to get the source files of all shortcuts (Apple, 2024b). We can share shortcuts using the Shortcuts app on Apple devices. The first method is through iCloud links, which do not provide access to the shortcut’s source file. The second method involves sharing the shortcut’s source file, with the .shortcut suffix, but these files are signed by Apple devices and cannot be easily decrypted. Through our efforts, we discovered that when shortcuts are imported into the Shortcuts app, they are displayed in an easily understandable PLIST file format. Users can obtain the source files of any shortcut using our decryption shortcuts available at https://www.icloud.com/shortcuts/b04412850b9f4f74ad16f2f15ef09a3f and https://www.icloud.com/shortcuts/8fa07dea82cf413c81732dca5f15323f. To facilitate the large-scale acquisition of shortcut source files, we analyzed the network traffic of the Shortcuts app. We found that we could retrieve shortcut metadata from https://www.icloud.com/shortcuts/api/records/\${unique_id}, where \${unique_id} is a string like b04412850b9f4f74ad16f2f15ef09a3f. Using the download link indicated by the metadata at ["fields"]["shortcut"]["value"]["downloadURL"], we could obtain the shortcut source files. This method enabled us to acquire all shortcut source files. The retrieved metadata includes the downloadURL and the name field of the shortcut. In our subsequent processing, we use this name to replace NameInStore, as it directly corresponds to the name of the shortcut when imported into the Shortcuts app.
Subsequently, we extracted “app name” using the  field WFWorkflowActionIdentifier in the shortcut source file like com.openai.chat.AskIntent, and then downloaded related apps from various sources. Shortcuts are composed of a series of shortcut API calls (Actions). A typical shortcut is shown in Figure 7. Each shortcut API call is identified by a name, which typically consists of the name of the shortcuts app like com.openai.chat and the name of the Intent like AskIntent. For most API names, the part before the last dot is the app name, and the part after is the Intent name. We semi-automatically extracted all app names to facilitate downloading the apps.
@@
Most app downloads can be accomplished using the ipatool tool, which supports downloading iOS / iPadOS versions of apps. However, there are two categories of apps that ipatool cannot download: (1) Some apps are available only in macOS and do not have iOS / iPadOS versions; (2) Most first-party Apple apps are not available for download from the App Store, and even those that do not have API definition files. For category (1), we manually downloaded these apps. For category (2), we found most of the apps in the /Applications/ and /System/Applications/ directories on macOS, and located most of the API definition files in /System/Library/PrivateFrameworks/WorkflowKit.framework/. In summary, we adopt the following approach to choose apps: (1) We select the macOS version for apps from Apple like the “Shortcuts” app. (2) For apps that have an iOS version, we prioritize the iOS version. (3) For apps that only have a macOS version, we choose the macOS version. Then we conducted app and shortcut filtering to remove apps that had no definition file and shortcuts that had APIs with missed API descriptions. Our choice here is made solely for the convenience of extracting API definition files. For most apps, selecting iOS / iPadOS or macOS versions does not impact the retrieval of API definition files. We found that only a few apps, such as Chrome, have different APIs on macOS compared to iOS / iPadOS. These differences will be filtered out in subsequent steps.
@@
Then we managed to extract APIs from the downloaded apps. The APIs are mainly from intent definition file \${filename}.actionsdata from AppIntent (Apple-Inc., 2024b; 2022; 2023) framework, \${filename}.intentdefinition from SiriKit (Apple-Inc., 2024e; 2022; 2023; c) framework, and WFActions.json from system path /System/Library/PrivateFrameworks/WorkflowKit.framework/ on macOS. SiriKit was introduced in 2016 with iOS 10 (Apple-Inc., 2024e), allowing applications to integrate with Siri for voice command interactions. AppIntents, launched with iOS 16 in 2022 (Apple-Inc., 2024b), offers a more modern and flexible way to define and handle app intents, facilitating integration with Siri, Shortcuts, widgets, and more. Apple is actively encouraging developers to adopt AppIntents by providing migration tools from SiriKit (Apple-Inc., 2023; 2024d). However, some apps still use the SiriKit framework. When developing with the SiriKit framework, \${filename}.intentdefinition files are used, while the App Intent framework uses \${filename}.actionsdata files. These files define the APIs corresponding to actions in shortcuts. An app may contain only \${filename}.intentdefinition files, only \${filename}.actionsdata files, or both, leading to potential redundancy in API definitions. We strive to reduce the number of API definition files and have established a set of rules to ensure API uniqueness in ShortcutsBench. Finally, as shown in Table 1, we get $88$ apps from various categories such as “Health & Fitness”(iTunes App Store, 2024b), “Developer Tools” (iTunes App Store, 2024a), and “Lifestyle” (iTunes App Store, 2024c). These apps in total include $1414$ APIs, including all of $556$ APIs involved in $7627$ shortcuts.
@@ A.2 Dataset Construction
We construct queries based on existing action sequences and APIs. To ensure the quality of these queries, we utilize the natural language workflow descriptions unique to shortcuts. When generating queries, we require the model to naturally include primitive data type parameters and enum data types needed for API calls. This helps us evaluate the agent’s ability to handle primitive parameters. We do not require the inclusion of complex data types in the queries, as they are difficult to convert to text and challenging to evaluate. To ensure high-quality query generation, we use the state-of-the-art LLM, GPT-4o (OpenAI, 2024b). The prompt templates used for generating queries are provided in Figure 9.
@@ A.4 Performance about API Selection
Following existing work (Huang et al., 2023; Patil et al., 2023; Li et al., 2023b; Xu et al., 2023; Schick et al., 2024; Hao et al., 2024), we use the accuracy of API selection as the metric. The accuracy is calculated as the number of correct API selections $m_{p}$ divided by $n_{p}$. Specifically, each time we predict an action $b_{j},1\leq j\leq|aseq_{i}|$, we provide the agent with all the correct historical actions $\{a_{1},a_{2},...,a_{j-1}\}$. We then require the agent to predict the next action. All actions predicted by the agent form the prediction sequence $bseq_{p,i}$. This method is similar to the next token prediction (NTP) in LLMs, effectively preventing a cascade of errors in subsequent action predictions due to a single incorrect prediction. During the prediction, when encountering special actions such as branching and looping, we skip predicting these actions and directly add them to the historical actions.
@@ A.6 Recognition of Need for Input
In the shortcut, a parameter can be set to ExtensionInput, indicating that the parameter requires a file provided by the user, or CurrentDate, indicating that the parameter needs to retrieve the date from the system. Similarly, Clipboard indicates that the parameter should obtain content from the clipboard, and DeviceDetails implies that the parameter needs to access certain information about the user’s device. Lastly, Ask denotes that the parameter requires user authorization or essential information from the user. A typical example is shown in Figure 12, where the action uses the is.workflow.actions.getmyworkflows API. The Folder parameter is set to Ask, indicating that this parameter requires information provided by the user.
@@ A.7 Setup
Following existing work (Huang et al., 2023; Qin et al., 2023; Patil et al., 2023; Tang et al., 2023; Li et al., 2023b; Xu et al., 2023; Zhuang et al., 2024; Schick et al., 2024; Hao et al., 2024), we slightly modified the ReACT (Yao et al., 2022) templates to construct the API-based agents. The templates used in our experiments are as shown in Figure 13.
@@ A.8 Result Analysis
Total Cost 3153
Among them, gemini-1.5-pro (tested with 801 instances) and gemini-1.5-flash (tested with 5,295 instances) incurred a total cost of \$801, with gemini-1.5-flash accounting for approximately \$391 and gemini-1.5-pro approximately \$592. The costs for qwen2-72b-instruct (tested with 5,216 instances) were about \$800, qwen2-57b-a14b-instruct (tested with 5,368 instances) around \$580, gpt-3.5-turbo (tested with 5,463 instances) approximately \$500, and the combined expenses for deepseek-chat (tested with 5,319 instances) and deepseek-coder (tested with 5,317 instances) were roughly \$180. GLM-4-Air cost about \$110.

### v3

@@ 5 Conclusion
In this paper, we introduce ShortcutsBench, a benchmark for evaluating API-based agents. To the best of our knowledge, ShortcutsBench is the most realistic, rich, and comprehensive benchmark of its kind. Our findings indicate that for agents built on the most advanced LLMs, the primary bottleneck is API selection. For the most cost-effective LLMs, there is considerable room for improvement in both API selection and parameter filling. Additionally, we identified a significant deficiency in the agents’ awareness of requesting necessary input.
@@ Appendix A Appendix / supplemental material
A.1 Dataset Acquisition
In this section, we introduce more details about the dataset acquisition introduced in Section 3.1.
Regarding data acquisition, we first use search engines to identify popular public shortcut-sharing sites (① in Figure 1). We found a total of $14$ sites. These sites include:
Table 5: 14 shortcut-sharing sites, including names, URLs, and shortcut counts.
@@ A.1 Dataset Acquisition
We can obtain shortcuts from these sites. Specifically, each dataset includes the “shortcut name” (NameInStore), “function description” (DescriptionInStore), “shortcut type” (CategoryInStore), and most importantly, the “iCloud link” (Apple, 2024a). Additionally, it includes less important data such as the number of downloads (Downloads), favorites (Favorites), reads (Reads), and ratings (Rates). All shortcuts include NameInStore and DescriptionInStore, while the availability of other fields varies slightly depending on the specific shortcut-sharing site.
We then downloaded the shortcut source file by “iCloud link” and performed deduplication based on both iCloud links and the actual shortcut content (i.e., action sequences) to ensure the uniqueness of each shortcut in the final dataset (② in Figure 1). For details on downloading source files via iCloud links, please refer to our open-source code repository. We do deduplication because shortcuts sharing sites store shortcuts as iCloud links, which often results in the same shortcut appearing in multiple sharing-site. Additionally, shortcuts linked by these iCloud links could have identical content, making deduplication essential to ensure that each shortcut in the final dataset was unique.
We then extracted the app name using the field WFWorkflowActionIdentifier from the shortcut source file and downloaded the associated apps (③ in Figure 1). Shortcuts are composed of a series of shortcut API calls, referred to as Actions. An example of a typical shortcut is shown in Figure 7. Each shortcut API call is identified by a name, which usually includes the app’s identifier, such as com.openai.chat, and the Intent name, such as AskIntent. For most API names, the segment before the last dot represents the app name, while the segment after denotes the Intent name. We semi-automatically extracted all app names to streamline the app download process.
@@
We download these apps from various sources:
  • Apps from the macOS or iOS App Store: We downloaded a variety of applications directly from Apple’s official platforms. This provided us with a vast selection of apps that are widely used and trusted by users.
  • System apps like Keynote from paths /Applications/ and /System/Application/ on macOS: These are pre-installed applications integral to the operating system. Including them ensured that our dataset covered essential tools commonly used by macOS users.
  • Third-party apps from the official websites of the apps: To include software not available through the App Store, we downloaded apps from their official websites. This allowed us to capture a broader range of functionalities offered by third-party developers.
During the downloading process, we also excluded some legacy apps that are no longer maintained and $12$ paid apps to avoid potential licensing issues and focus on applications readily accessible to the general public.
Then we managed to extract APIs from the downloaded apps (④ in Figure 1). The APIs are mainly from intent definition file \${filename}.actionsdata from AppIntent (Apple-Inc., 2024b) framework and \${filename}.intentdefinition from SiriKit (Apple-Inc., 2024c) framework. We extracted all APIs involved in the apps. During the extraction, we perform deduplication of APIs based on manually crafted rules as an app may have multiple duplicate API definition files with the same API definition.
@@
We perform deduplication to streamline API definitions, minimize redundancy, and ensure compatibility across frameworks, addressing inconsistencies introduced by the coexistence of SiriKit and AppIntents. SiriKit, introduced in 2016 with iOS 10, enabled applications to integrate with Siri for voice command interactions. In 2022, Apple launched AppIntents with iOS 16, providing a more modern and flexible approach to defining and handling app intents. AppIntents facilitate integration with Siri, Shortcuts, widgets, and more. To encourage adoption, Apple has provided migration tools for developers transitioning from SiriKit. However, some apps still rely on the SiriKit. Under SiriKit, developers use \$filename.intentdefinition files, while the AppIntents relies on \$filename.actionsdata files. These files define APIs corresponding to actions in Shortcuts. Apps may include only \$filename.intentdefinition files, only \$filename.actionsdata files, or both, potentially leading to redundancy in API definitions. To address this, we have implemented a set of rules to reduce API definition files and ensure API uniqueness.
Additionally, for app Shortcuts, which are deeply integrated with Apple’s operating system, we need to obtain their API definition files WFActions.json from system path /System/Library/PrivateFrameworks/WorkflowKit.framework/ on macOS, instead of extracting it from the app itself.
Subsequently, we further filtered the shortcuts based on criteria such as whether the associated apps is paid app, whether the apps were outdated, and whether the APIs were deprecated. Additionally, we imported all shortcuts into the macOS Shortcuts app to ensure they were functional. These steps were repeated multiple times.
Finally, as shown in Table 2, we get $88$ apps from various categories such as “Health & Fitness”, “Developer Tools”, and “Lifestyle”. These apps in total include $1414$ APIs, including all of $556$ APIs (Not all APIs have been used in Shortcuts) involved in $7627$ shortcuts.
The approximate time spent on each step of the process is outlined below:
  • Shortcut site collection: Approximately 3 days, completed entirely manually.
  • Link scraping using Selenium: Around 2 weeks, requiring custom scripts for each site.
  • Shortcut deduplication, API validity checks, and shortcut functionality validation: Approximately 4 weeks. Deduplication: Automated using iCloud links and content cleaning.
  • API validity checks: Performed manually. Shortcut validity checks: A mix of automated and semi-automated methods. Automated filtering was conducted using Apple Scripts to execute shortcuts for preliminary filtering, followed by manual validation through importing shortcuts into the Shortcuts app.
Additional manual and automated checks were conducted throughout the process but are not detailed here.
@@ A.2 Dataset Construction
We construct queries based on existing action sequences and APIs. To ensure the quality of these queries, we utilize the natural language workflow descriptions unique to shortcuts. When generating queries, we require the model to naturally include primitive data type parameters and enum data types needed for API calls. This helps us evaluate the agent’s ability to handle primitive parameters. We do not require the inclusion of complex data types in the queries, as they are difficult to convert to text and challenging to evaluate. To ensure high-quality query generation, we use the state-of-the-art LLM, GPT-4o (OpenAI, 2024). The prompt templates used for generating queries are provided in Figure 9.
To ensure the quality of shortcuts, we filtered them based on criteria such as whether the associated apps were paid, outdated, or relied on deprecated APIs. Additionally, all shortcuts were imported into the macOS Shortcuts app to verify functionality. Deduplication and error-checking processes were carried out throughout the entire data collection phase.
For ensuring the quality of generated queries, following prior work (Qin et al., 2024), we conducted a preliminary experiment with three LLMs: GPT-4o, GPT-3.5, and Gemini-1.5-Pro, on a dataset of 100 samples. Human evaluators rated GPT-4o as generating the highest-quality queries, outperforming the other two models. GPT-4o excelled in accurately identifying required parameters and providing clear query descriptions, meeting our criteria in 94 out of 100 cases. This superior performance can largely be attributed to the natural language workflow descriptions. While we acknowledge that not all queries may fully meet our requirements, we believe our approach is reasonable. Similar works, such as ToolLLM, rely on GPT for large-scale query and action sequence generation without guaranteeing complete accuracy.
@@ A.4 Performance about API Selection
Following existing work (Huang et al., 2024b; Patil et al., 2024; Xu et al., 2024), we use the accuracy of API selection as the metric. The accuracy is calculated as the number of correct API selections $m_{p}$ divided by $n_{p}$. Specifically, each time we predict an action $b_{j},1\leq j\leq|aseq_{i}|$, we provide the agent with all the correct historical actions $\{a_{1},a_{2},...,a_{j-1}\}$. We then require the agent to predict the next action. All actions predicted by the agent form the prediction sequence $bseq_{p,i}$. This method is similar to the next token prediction (NTP) in LLMs, effectively preventing a cascade of errors in subsequent action predictions due to a single incorrect prediction. During the prediction, when encountering special actions such as branching and looping, we skip predicting these actions and directly add them to the historical actions.
@@ A.6 Recognition of Need for Input
In the shortcut, a parameter can be set to ExtensionInput, indicating that the parameter requires a file provided by the user, or CurrentDate, indicating that the parameter needs to retrieve the date from the system. Similarly, Clipboard indicates that the parameter should obtain content from the clipboard, and DeviceDetails implies that the parameter needs to access certain information about the user’s device. Lastly, Ask denotes that the parameter requires user authorization or essential input from the user. A typical example is shown in Figure 12, where the action uses the is.workflow.actions.getmyworkflows API. The Folder parameter is set to Ask, indicating that this parameter requires input provided by the user.
@@ A.7 Setup
Following existing work (Huang et al., 2024b; Qin et al., 2024; Li et al., 2023), we slightly modified the ReACT (Yao et al., 2023) templates to construct the API-based agents. The templates used in our experiments are as shown in Figure 13.
@@ A.8 Result Analysis
Total Cost 3253
Among them, gemini-1.5-pro (tested with 801 instances) and gemini-1.5-flash (tested with 5,295 instances) incurred a total cost of \$801, with gemini-1.5-flash accounting for approximately \$391 and gemini-1.5-pro approximately \$592. The costs for qwen2-72b-instruct (tested with 5,216 instances) were about \$800, qwen2-57b-a14b-instruct (tested with 5,368 instances) around \$580, and GPT-4o-mini (tested with 5,320 instances) approximately \$50. gpt-3.5-turbo (tested with 5,463 instances) cost approximately \$500. The combined expenses for deepseek-chat (tested with 5,319 instances) and deepseek-coder (tested with 5,317 instances) were roughly \$180, while GLM-4-Air cost about \$110.
@@
A.9 Additional Results on Evaluation on specialized API-calling fine-tuned LLMs
To comprehensively evaluate the effectiveness of fine-tuning LLMs for agent-based scenarios, we extended our comparison to include specialized AI agent models such as AgentLM (Zeng et al., 2024) and xLAM (Zhang et al., 2024), alongside general-purpose LLMs like Qwen-2-72B Qwen (2024) and LLaMA-3-70B Meta (2024). This was aimed at understanding the applicability of fine-tuning strategies for API-based agents.
We conducted experiments with additional $11$ LLMs, comprising $6$ models fine-tuned specifically for agent scenarios and $5$ base models. The fine-tuned models include the AgentLM series (AgentLM-7B, AgentLM-13B, AgentLM-70B), the xLAM series (xLAM-7b-r, xLAM-8x7b-r), and Lemur (Lemur-70B-Chat-V1). The base models include various configurations of LLaMA-2 and Mistral from Hugging Face repositories.
The fine-tuning processes of these LLMs varied: the AgentLM series was fine-tuned on $1866$ traces of simple code data, Lemur utilized scripting and interpreted languages from The Stack dataset, and the xLAM series was specialized for API-based agents with a predefined format for API calls. To ensure standardized evaluation, outputs were processed using gpt-4o-mini, with manual verification confirming the accuracy of conversions. Due to context length limitations (4k tokens for most models and 32k tokens for xLAM series), results are reported for the first $3$ difficulty levels.
Experimental results are summarized in Table 7.
Table 7: Performance Comparison of Fine-Tuned and Base LLMs
Model
LLaMA-2-7B
AgentLM-7B
LLaMA-2-13B
AgentLM-13B
LLaMA-2-70B
AgentLM-70B
Base Fine-Tuned Base Fine-Tuned Base Fine-Tuned
Context Length 4k 4k 4k 4k 4k 4k
API Selection 32.75 24.95 38.59 30.57 52.13 43.41
12.57 17.43 25.99 35.48 43.08 37.34
23.70 32.78 35.45 40.78 57.76 39.22
API Parameter Value Filling 51.11 34.88 30.77 31.82 72.86 52.83
32.14 35.71 12.50 40.98 45.05 37.31
0.00 0.00 11.11 40.00 6.67 11.11
Recognition of Need for Input 7.41 10.53 5.00 5.26 21.74 0.00
12.50 4.17 0.00 3.57 7.69 9.09
0.00 0.00 0.00 0.00 0.00 0.00
Model
Mistral-7B
xLAM-7b-r
Mixtral-8x7b
xLAM-8x7b-r
LLaMA-2-70B
Lemur-70B-Chat-V1
Base Fine-Tuned Base Fine-Tuned Base Fine-Tuned
Context Length 32k 32k 32k 32k 4k 4k
API Selection 17.03 70.56 70.21 85.29 52.13 52.45
3.27 50.82 49.15 51.72 43.08 47.08
0.50 66.33 29.92 29.67 57.76 60.39
API Parameter Value Filling 40.00 66.67 53.45 68.93 72.86 45.45
40.54 59.32 33.72 46.76 45.05 36.49
11.11 31.82 27.20 42.71 6.67 26.92
Recognition of Need for Input 0.00 5.56 6.25 11.90 21.74 16.00
0.00 2.56 19.27 19.91 7.69 3.57
0.00 0.00 0.00 0.00 0.00 0.00
A.9.1 Recognition of Need for Input
All evaluated LLMs performed exceptionally poorly in the recognition of need for input, significantly lagging behind the general LLMs mentioned in the ShortcutsBench. Compared to the performance of the base LLMs before fine-tuning, neither the AgentLM and Lemur series models fine-tuned on code nor the xLAM series models fine-tuned specifically for API-based agents showed any improvement in this aspect. This suggests the need for targeted fine-tuning specifically aimed at enhancing this capability.
A.9.2 API Parameter Value Filling
The evaluated LLMs perform significantly worse in API Parameter Value Filling compared to the general LLMs mentioned in ShortcutsBench. The general LLMs achieve accuracies of 94.29%, 83.92%, and 75.53% across three difficulty levels, whereas the best results from the evaluated models only reach accuracies of 72.86%, 59.32%, and 42.71%.
Compared to the base LLMs before fine-tuning, the code-based agents, including the AgentLM and Lemur models fine-tuned on code, did not achieve a general improvement in parameter-filling accuracy for API-based tasks and even exhibited performance degradation. In contrast, the API-based agents, such as the xLAM series fine-tuned specifically for API-based agents, showed a consistent improvement in parameter-filling accuracy. The existing fine-tuning methods for API-based agents (xLAM-7b-r / xLAM-8x7b-r) are effective for parameter filling in short sequences, but accuracy drops sharply as the context length increases.
A.9.3 API Selection
Methods fine-tuned specifically for API-based agents, such as the xLAM series models, can significantly enhance agents’ API selection accuracy. The xLAM series achieves reasonable accuracy in API selection, even compared to general LLMs. However, compared to xLAM-7b-r, the API selection accuracy of xLAM-8x7b-r is significantly lower, suggesting that existing fine-tuning methods specialized for agents may only achieve notable improvements on models with relatively weaker foundational capabilities.
Additionally, methods fine-tuned on code can still enhance agents’ API selection accuracy, as demonstrated by the AgentLM-7b/13b and Lemur-70b-chat-v1 models. The performance drop of AgentLM-70b may be attributed to the fine-tuning dataset used for the AgentLM series lacking realism and complexity.